help me please for the solution,,,,,,, three angles of a triangle measure x+10, 2x+15, and 3x-13. what are the degree measures of the angles of the triangle?
Answer by mananth(16949)

x+10, 2x+15, and 3x-13. what are the degree measures of the angles of the triangle?
The sum of the measures of the angles of triangle = 180 degrees
x+10+2x+15+3x-13=180
6x+12=180
-12 -12
6x=168
/6
x=28
plug x in the angles and you will get the measures
x+10 = 28+10=38
2x+16 = 2*28+15=71deg
you work out the third